OpenStudy (anonymous):

CeeCee recorded the number of necklaces for sale at 10 different locations and then arranged the numbers from smallest to largest. The numbers of necklaces for sale at the locations were 4, 4, 4, 5, 8, 9, 9, 9, 9, and 9.The mean number of necklaces for sale at the locations was 7. What is the mean absolute deviation? A. 0 B. 2 C. 2.2 D. 4.4

OpenStudy (anonymous):

pleas the last day

OpenStudy (anonymous):

add all the numbers it gives you and divide by 10 then subtract whatever you get here ^ with each number. add all these numbers ^ divide by 10 answer^

OpenStudy (anonymous):

what

OpenStudy (anonymous):

7

OpenStudy (texaschic101):

take each number on your data set and subtract the mean from it....and even if it is negative, mark it positive 4 - 7 = 3 4 - 7 = 3 4 - 7 = 3 5 - 7 = 2 8 - 7 = 1 9 - 7 = 2 9 - 7 = 2 9 - 7 = 2 9 - 7 = 2 9 - 7 = 2 now find the mean of that (3 + 3 + 3 + 2 + 1 + 2 + 2 + 2 + 2 + 2) / 10 = your answer

OpenStudy (anonymous):

2.2

OpenStudy (texaschic101):

correct :)

OpenStudy (anonymous):

okay thanks can you help me with 2 more pleas

OpenStudy (texaschic101):

I can try...can't make no promises

OpenStudy (anonymous):

okay thanks

OpenStudy (anonymous):

Uri asked six friends how many different foreign countries they’d been to. He recorded 0, 1, 1, 2, 5, and 9 countries visited. What is the mean absolute deviation of the number of countries visited? Round to the nearest tenth, if needed. A. 0 countries B. 2.7 countries C. 2.9 countries D. 3 countries I think it A

OpenStudy (texaschic101):

first we find the mean of the original data (0 + 1 + 1 + 2 + 5 + 9) / 6 = 18/6 = 3 now subtract the mean from every number in the data set...making the answer positive 0 - 3 = 3 1 - 3 = 2 1 - 3 = 2 2 - 3 = 1 5 - 3 = 2 9 - 3 = 6 now find the mean of those numbers (3 + 2 + 2 + 1 + 2 + 6)/6 = 16/6 = 2.6666 rounds to 2.7 its B questions ?

OpenStudy (anonymous):

okay one more

OpenStudy (texaschic101):

I am ready :)

OpenStudy (anonymous):

Four friends played mini golf. Their scores were –3, –2, –1, and 2. What is the mean absolute deviation of their scores?

OpenStudy (anonymous):

I think its 1.5

OpenStudy (anonymous):

what do you think

OpenStudy (texaschic101):

(-3 -2 -1 + 2) / 4 = (-6 + 2)/4 = -1 now subtract mean from data set....making answers positive -3 - (-1) = -3 + 1 = 2 -2 - (-1) = -2 + 1 = 1 -1 - (-1) = -1 + 1 = 0 2 - (-1) = 2 + 1 = 3 (2 + 1 + 0 + 3) / 4 = 6/4 = 1.5 I think it is 1.5...but not 100% sure....the negatives are kinda throwing me off

OpenStudy (anonymous):

okay thanks let me submit it and i will tell you

OpenStudy (texaschic101):

sounds good :)

OpenStudy (anonymous):

yesss I got 80 % thanks so much

OpenStudy (texaschic101):

sweet :) good job...glad to help